# IDEA: Enhancing the Rule Learning Ability of Large Language Model Agent through Induction, Deduction, and Abduction

### v5

@@ Abstract
While large language models (LLMs) have been thoroughly evaluated for deductive and inductive reasoning, their proficiency in holistic rule learning in interactive environments remains less explored. We introduce RULEARN, a novel benchmark to assess the rule-learning abilities of LLM agents in interactive settings. In RULEARN, agents strategically interact with simulated environments to gather observations, discern patterns, and solve complex problems. To enhance the rule-learning capabilities for LLM agents, we propose IDEA, a novel reasoning framework that integrates the process of Induction, DEduction, and Abduction. The IDEA agent generates initial hypotheses from limited observations through abduction, devises plans to validate these hypotheses or leverages them to solve problems via deduction, and refines previous hypotheses through induction, dynamically establishing and applying rules that mimic human rule-learning behaviors. Our evaluation of the IDEA framework, which involves five representative LLMs, demonstrates significant improvements over the baseline. Furthermore, our study with human participants reveals notable discrepancies in rule-learning behaviors between humans and LLMs. We believe our benchmark will serve as a valuable and challenging resource, and IDEA will provide crucial insights for the development of LLM agents capable of human-like rule learning in real-world scenarios. Our code and data is publicly available [^1] https://github.com/KaiyuHe998/RULEARN_IDEA.
@@ 1 Introduction
One major pillar of human intelligence is the ability to discern rules and apply them. We identify patterns, formulate hypotheses, and refine them by interacting with the environment. This exploratory process traditionally involves three stages: abduction, deduction, and induction. According to Charles Peirce’s definition (Frankfurt, 1958; Peirce, 1974), the rule-learning loop typically begins with an explanatory hypothesis that arises from abduction. This is followed by iterative experiments guided by the hypothesis, known as deduction, which leads to the modification and refinement of the hypothesis through induction (As shown in Figure 1). This establishes the foundation of how humans learn the rules of the world and apply them to effect change. In the real world, such a process involves iteratively refining hypotheses through interaction and evidence gathering, driven by a dynamic interplay between reasoning stages and the environment. However, recent studies often isolate abduction, induction, and deduction in non-interactive settings, missing this interdependent nature (Bowen et al., 2024; Wang et al., 2023; Saparov et al., 2024; Liu et al., 2024).
@@
To address this gap, we introduce RULEARN, a new benchmark to evaluate the rule-learning abilities of LLM agents in interactive environments. RULEARN consists of 300 high-quality puzzles with hidden rules set in a text-based environment, where agents begin exploration without prior knowledge of these rules. Successfully solving the puzzles requires the agent to strategically select actions, efficiently gather pattern-revealing observations, and accurately reason from them to infer the hidden rules. RULEARN presents substantial challenges, as agents must rely entirely on observations generated by their chosen actions to discern rules. If these observations fail to reveal a clear pattern, the agent is likely to fail.
We design three types of environments within RULEARN to evaluate the rule-learning ability in different scenarios: (1) The Function Operator: Determining the coefficients of mathematical functions defined by hidden expressions. Agents can assign various values to the input variables and observe the outputs, using this information to hypothesize the function’s form. The challenge lies in efficiently selecting input values that reveal the underlying structure and accurately computing the coefficients based on limited observations. (2) The Escape Room: Deciphering the passcode to exit an escape room. A hidden rule determines how the objects in the room infer the passcode digits. Agents interact with these objects to gather clues and input passcodes into the door. Based on feedback, agents formulate hypotheses and infer the relationship between the objects and the passcode. (3) The Reactor: Synthesizing target strings using a reactor with a hidden string-combining rule. Agents need to experiment with different inputs and analyze outputs to deduce the reactor’s transformation rule and achieve the desired outcome.
@@
We evaluate IDEA on five popular LLMs—GPT-3.5-Turbo, GPT-4o, Gemma-7B, Llama3-8B, and Llama3-70B—observing roughly a 10% improvement in success rates compared to the baseline. Without hypothesis guidance, the baseline agent tends to choose direct, uninformed actions that fail to uncover the hidden rules. In contrast, the IDEA agent reduces repeated actions by 30.2%, obtains more diverse observations, and better understands the underlying rules. To further investigate their rule-learning capabilities, we compare LLM performance with that of 50 human participants. Although IDEA narrows the gap, LLMs still face challenges: (1) inefficient exploration in unfamiliar environments, resulting in insufficient evidence to reveal rules; (2) difficulty in deducing valid plans to verify current hypotheses and guide future exploration; and (3) reluctance to correct initial hypotheses when confronted with contradictory observations. These findings provide important insights into improving LLM agents to achieve more human-like rule learning in complex, real-world scenarios.
@@ 2 Related Works
Agents powered by large language models (LLMs) with memory, tool use, planning, and reasoning have shown notable progress in understanding complex tasks (Wang et al., 2024b; Chen et al., 2023; Zhou et al., 2023; Wang et al., 2024a; Andreas, 2022; Park et al., 2023; Zhong et al., 2023; Zhang et al., 2024b; Nakano et al., 2022; Lu et al., 2023; Shi et al., 2023a; Schick et al., 2023; Yuan et al., 2023; Shen et al., 2023; Yao et al., 2023a; Besta et al., 2024). Recent work examines different reasoning processes (abduction, deduction, induction) in LLMs (Bowen et al., 2024; Wang et al., 2023; Saparov et al., 2024; Cheng et al., 2024; Yang et al., 2024), but typically in isolation. As a result, their comprehensive rule-learning abilities in interactive settings remain underexplored.
Current reasoning tasks are hindered by inadequate benchmarks, primarily relying on QA datasets like Hotpot-QA (Yang et al., 2018) and Trivia-QA (Joshi et al., 2017) that lack the requirement for active information gathering, and coarse-grained interactive environments such as TextWorld (Côté et al., 2019) and AlfWorld (Shridhar et al., 2021) that limit agents to high-level actions (e.g., go to, open), impeding complex, experiment-driven rule discovery. This approach contrasts with real-world rule learning, which necessitates active evidence gathering, experimentation, and iterative refinement. Exacerbating this issue, many existing studies employ static, non-interactive scenarios where LLMs receive all information upfront (Yang et al., 2023; Zhu et al., 2024; Shi et al., 2023b; Liu et al., 2024), failing to mirror the dynamic nature of real-world learning processes. Furthermore, while recent efforts (Xu et al., 2024; Montes et al., 2022) attempt to integrate interactivity, they still treat information gathering, rule generation, and application as distinct phases, thereby undermining the development of agents capable of seamless integration of these elements in real-world settings.
This gap arises due to the absence of an interactive, text-based environment that supports fine-grained actions for rule learning. To address this, RULEARN enables actions where agents can input arbitrary strings, and our system parses these inputs, providing tailored feedback on a character-by-character basis. This capability allows LLM agents to conduct fine-grained experiments and closely mirror real-world rule learning. In turn, IDEA equips agents to effectively manage the interdependent processes of information gathering, hypothesis generation, and validation in a unified manner, offering a more comprehensive framework that mimics human rule-learning behaviors.
3 The RULEARN Benchmark
We develop three puzzle sets—Function Operator, Escape Room, and Reactor—each consisting of 100 unique puzzles of varying complexity, all manually created by our coauthors, who are computer science graduate students. Unlike existing fine-grained interactive environments, which are predominantly found in the robotics domain and offer significantly fewer tasks (Jain et al., 2020; Nasiriany et al., 2024; Zhang et al., 2024a), RULEARN is the first text-based environment providing such fine-grained interactions specifically for language agents.
@@ 3 The RULEARN Benchmark
The Function Operator. This puzzle type simulates scenarios where systemic theories or established knowledge (in this case, mathematics) is applicable. The agent interacts with a set of univariate multi-term equations involving integer parameters from [0,9] and elementary functions of the variable $x$, selected from $f(x)\in\{x^{0},x^{1},x^{2},\sin(x),\frac{1}{x},|x|,-x\}$. The agent is provided with the number of functions, the presence of specific parameters in each function (the exact numerical values of these parameters are unknown and represented by letters), and the types of elementary functions involved in the current puzzle. The goal of the agent is to deduce the values of these parameters. For example, in one puzzle, the ground truth is $\mathbf{F_{1}}(x)=a\sin(x)+b\times\frac{1}{x}$, $\mathbf{F_{2}}(x)=ax^{2}$ where $a=3$ and $b=2$. The agent knows the following information: There are three elementary functions in this puzzle $\{\sin(x),\frac{1}{x},x^{2}\}$, there are two functions $\mathbf{F_{1}}(x)$ and $\mathbf{F_{2}}(x)$, $\mathbf{F_{1}}(x)$ has 2 terms and parameters $a,b$ in it, and $\mathbf{F_{2}}(x)$ has 1 term and one parameter $a$. To solve the puzzle, the agent must interact with the environment through a defined action space: selecting a function and assigning values to $x$, then observing the resultant output. For example, assigning values 1 and 2 to $\mathbf{F_{2}}$ reveals a quadratic increase in output, indicating the presence of $x^{2}$ in $\mathbf{F_{2}}$. Similarly, assigning a value of 1 to $\mathbf{F_{1}}$ results in a floating-point output, rather than an integer, suggesting the inclusion of trigonometric components, confirming that $\sin(x)$ is a component of $\mathbf{F_{1}}$. The difficulty of each puzzle is controlled by variations in the number of functions, unknown parameters, and elementary functions in use. We manually enumerated 100 combinations of functions, incorporating different numbers of terms and types of elementary functions to ensure a diverse range of puzzle complexity (see detailed distribution in Table 4 in Appendix  A.2).
@@
The Escape Room. This environment simulates scenarios where no established knowledge is applicable. We create a fictitious setting: an agent is placed in an art gallery escape room and must decipher a 3-digit password to unlock a code-secured door. Each digit of the password represents the count of paintings of a specific type—watercolor, oil, or acrylic—that share a given color. The agent receives brief descriptions of paintings, such as “This is an acrylic painting of a green jungle”, indicating their type and color. Initially, the agent only knows the password is a 3-digit number and is given a hint about which color to focus on. After proposing a hypothesis and entering a password guess, the door provides feedback on which digits are correct, allowing the agent to refine its hypothesis. To prevent a brute-force approach, the specific color associated with the password changes after every three failed attempts. Each puzzle varies in the number of paintings, and while paintings in the same room are visible, those in other rooms remain hidden until the agent moves to access them. The difficulty of this puzzle type is controlled by the different number of paintings, whether agent need to as detailed in Table 5 in the Appendix  A.2.
@@
The Reactor. This environment type also simulates scenarios where no established knowledge is applicable. However, unlike in the Escape Room where agents only input passwords to test hypotheses, in the Reactor, agents are challenged to design more fine-grained experiments to uncover hidden rules. Specifically, the agent’s task is to synthesize target strings using a reactor governed by a hidden string-combining rule. These strings are represented by sequences of alphabetic letters, such as A, B, AABB, and CAB. The reactor permits the agent to input two strings, initiating a reaction that produces a new string for use in subsequent experiments. The agent’s objective is to decipher the specific rules that govern string synthesis by methodically testing different string combinations, with the ultimate goal of synthesizing the target string using the discovered rules. We have designed four types of rules, detailed in Table 1. The difficulty of this puzzle type is controlled by the specific rules used, the length of the target string to be synthesized and the number of unique letters contained in the target string, as detailed in Table 6 in Appendix A.2.
@@
The detailed statistics for each puzzle type and example puzzles are provided in Appendix A.2 and A.7, respectively. Note that the RULEARN benchmark is designed to simulate real-world interactions and is adaptable to complex text environments with diverse rules. To achieve this, we do not restrict the representation of rules to a specific formal language, as doing so would limit realism and complexity. Moreover, predefining the rule format would provide agents with prior knowledge, contradicting real-world conditions. Therefore, we enable the LLM agents to use natural language to describe rules in RULEARN, promoting generalizability and future extensions of the benchmark.
@@ 4 The IDEA Agent
The IDEA agent begins with an abductive action to generate an initial hypothesis, followed by a deduction step to create a new plan. Based on this plan, the agent interacts with the environment. Upon receiving feedback from the environment as a new observation, the agent may take an inductive action to refine the hypothesis or perform another interaction with the environment. Deductive action is invoked to adjust the plan every time the hypothesis changes. This reasoning loop continues until the puzzle is solved or a maximum number of steps is reached. After each step, the results are appended to the agent’s memory, including interaction outcomes and any modifications to the hypothesis or plan. We provide a simplified algorithm demonstrating how the IDEA agent operates in Algorithm 1. Specifically, at each step, we prompt the LLM to reflect on the information recorded in the IDEA agent’s components to make decisions and take actions. We employ the chain-of-thought (CoT) reasoning (Wei et al., 2022) for all prompts. Detailed prompts for each type of action are available in Appendix A.5.
More detailed implementation of the agent can be found in Appendix A.3.1. Similar to real-life scenarios, when agents solve tasks in RULEARN puzzles, they do not know the outcomes in advance. Consequently, it is challenging to decide when to refine or change their hypothesis and plans, as well as what interactive actions to take to gather pattern-revealing observations. A detailed example of the IDEA agent solving the Reactor Puzzle is provided in Figure 5 in Appendix.

### v6

@@ Abstract
While large language models (LLMs) have been thoroughly evaluated for deductive and inductive reasoning, their proficiency in holistic rule learning in interactive environments remains less explored. We introduce RULEARN, a novel benchmark to assess the rule-learning abilities of LLM agents in interactive settings. In RULEARN, agents strategically interact with simulated environments to gather observations, discern patterns, and solve complex problems. To enhance the rule-learning capabilities for LLM agents, we propose IDEA, a novel reasoning framework that integrates the process of Induction, DEduction, and Abduction. The IDEA agent generates initial hypotheses from limited observations through abduction, devises plans to validate these hypotheses or leverages them to solve problems via deduction, and refines previous hypotheses through induction, dynamically establishing and applying rules that mimic human rule-learning behaviors. Our evaluation of the IDEA framework, which involves five representative LLMs, demonstrates significant improvements over the baseline. Furthermore, our study with human participants reveals notable discrepancies in rule-learning behaviors between humans and LLMs. We believe our benchmark will serve as a valuable and challenging resource, and IDEA will provide crucial insights for the development of LLM agents capable of human-like rule learning in real-world scenarios. Our code and data have been released on GitHub. [^1] https://github.com/KaiyuHe998/RULEARN_IDEA
@@ 1 Introduction
One major pillar of human intelligence is the ability to discern and apply rules. We identify patterns, form hypotheses, and iteratively refine them through interaction with the environment—a process that traditionally involves three stages: abduction, deduction, and induction. According to Charles Peirce (Frankfurt, 1958; Peirce, 1974), rule learning begins with an explanatory hypothesis (abduction), followed by iterative experiments (deduction) and hypothesis refinement (induction) (see Figure 1). This interdependent process underpins human rule learning in the real world, yet recent studies often isolate these stages in non-interactive settings (Bowen et al., 2024; Wang et al., 2023; Saparov et al., 2024; Liu et al., 2024).
@@
To simulate the full complexity of human rule learning, three essential principles must be met: an interactive environment that encourages dynamic experimentation, a fine-grained action space that enables flexible and precise inputs for nuanced hypothesis testing, and the presence of unknown rules that force agents to infer, test, and revise hypotheses based solely on sparse observations. By integrating these three key principles, we introduce RULEARN, which features 300 high-quality, manually created puzzles with hidden rules set in a text-based environment, where agents begin exploration without any prior knowledge of the underlying rules. RULEARN simulates human-like rule learning—compelling agents to experiment dynamically, utilize fine-graind actions for detailed hypothesis testing, and infer rules from limited data.
Successfully solving the puzzles requires the agent to strategically select actions, efficiently gather pattern-revealing observations, and accurately reason from them to infer the hidden rules. RULEARN presents substantial challenges, as agents must rely on observations generated by their chosen actions to discern rules; without clear patterns emerging from their exploration, the agents are likely to fail.
RULEARN consists of three types of environments to evaluate the rule-learning ability in different scenarios: (1) The Function Operator: Determining the coefficients of mathematical functions defined by hidden expressions. Agents can assign various values to the input variables and observe the outputs, using this information to hypothesize the function’s form. The challenge lies in efficiently selecting input values that reveal the underlying structure and accurately computing the coefficients based on limited observations. (2) The Escape Room: Deciphering the passcode to exit an escape room. A hidden rule determines how the objects in the room infer the passcode digits. Agents interact with these objects to gather clues and input passcodes into the door. Based on feedback, agents formulate hypotheses and infer the relationship between the objects and the passcode. (3) The Reactor: Synthesizing target strings using a reactor with a hidden string-combining rule. Agents need to experiment with different inputs and analyze outputs to deduce the reactor’s transformation rule and achieve the desired outcome.
@@
We evaluate IDEA on five popular LLMs—GPT-3.5-Turbo, GPT-4o, Gemma-7B, Llama3-8B, and Llama3-70B—observing roughly a 10% improvement in success rates compared to the baseline. Without hypothesis guidance, the baseline agent tends to choose direct, uninformed actions that fail to uncover the hidden rules. In contrast, the IDEA agent reduces repeated actions by 30.2%, obtains more diverse observations, and better understands the underlying rules. To further investigate their rule-learning capabilities, we compare LLM performance with that of 50 human participants. Although IDEA narrows the gap, LLMs still face challenges: (1) inefficient exploration in unfamiliar environments, resulting in insufficient evidence to reveal rules; (2) difficulty in deducing valid plans to verify current hypotheses and guide future exploration; and (3) reluctance to correct initial hypotheses when confronted with contradictory observations. These findings provide important insights into improving LLM agents to achieve more efficient rule learning in complex scenarios.
@@ 2 Related Works
Agents powered by large language models (LLMs) have shown notable progress in understanding complex tasks (Wang et al., 2024b; Chen et al., 2023; Zhou et al., 2023; Wang et al., 2024a; Andreas, 2022; Park et al., 2023; Zhong et al., 2023; Zhang et al., 2024b; Nakano et al., 2022; Lu et al., 2023; Shi et al., 2023a; Schick et al., 2023; Yuan et al., 2023; Shen et al., 2023; Yao et al., 2023a; Besta et al., 2024). Recent work examines different reasoning processes (abduction, deduction, induction) in LLMs (Bowen et al., 2024; Wang et al., 2023; Saparov et al., 2024; Cheng et al., 2024; Yang et al., 2024), but typically in isolation. As a result, their comprehensive rule-learning abilities in interactive settings remain underexplored.
@@
Current reasoning tasks are hindered by inadequate benchmarks that either rely on QA datasets like Hotpot-QA (Yang et al., 2018) and Trivia-QA (Joshi et al., 2017)—which lack active information gathering—or by coarse-grained interactive environments such as TextWorld (Côté et al., 2019) and AlfWorld (Shridhar et al., 2021) that limit agents to high-level actions (e.g., go to, open), impeding complex, experiment-driven rule discovery. This is in stark contrast to real-world rule learning, which requires active evidence gathering, experimentation, and iterative refinement. Moreover, many studies use static, non-interactive scenarios where LLMs receive all information upfront (Yang et al., 2023; Zhu et al., 2024; Shi et al., 2023b; Liu et al., 2024), failing to capture the dynamic nature of real-world learning. Even recent efforts (Xu et al., 2024; Montes et al., 2022) that integrate interactivity treat information gathering, rule generation, and application as distinct phases, undermining the development of agents capable of seamlessly integrating these elements.
We claim that to fully capture rule-learning ability in the real world, three criteria should be met: (i) Interactive Environment: the environment must be interactive so agents learn from interaction rather than passively receiving data; (ii) Fine-graind Action Space: the action space must be fine-grained, since existing benchmarks offer only coarse actions (e.g., go to, open) that prevent agents from performing the detailed experiments needed to test hypotheses (Jansen et al., 2024); and (iii) Unknown-Rule: the target rules must be unknown, as in ScienceWorld (Wang et al., 2022) most test cases rely on knowledge LLMs have already mastered during pretraining. To address this gap, our proposed RULEARN fulfills all three requirements by providing a fully interactive environment, offering a fine-grained action space in which agents submit arbitrary strings that our system parses to deliver character-level feedback, and introducing new, manually crafted puzzles whose rules are not familiar to LLMs. In turn, IDEA equips agents to manage the interdependent processes of information gathering, hypothesis generation, and validation within a unified framework that closely mimics human rule-learning behavior.
3 The RULEARN Benchmark
We develop three puzzle sets—Function Operator, Escape Room, and Reactor—each consisting of 100 unique, manually created puzzles of varying complexity, with each set reflects a different real-world rule-learning scenario. Unlike existing fine-grained interactive environments, which are predominantly found in the robotics domain and offer significantly fewer tasks (Jain et al., 2020; Nasiriany et al., 2024; Zhang et al., 2024a), RULEARN is the first text-based environment providing such fine-grained interactions specifically for language agents.
@@ 3 The RULEARN Benchmark
The Function Operator. This puzzle type simulates scenarios where systemic theories or established knowledge (e.g., mathematics) are applicable for efficiently testing hypotheses. The agent interacts with a set of univariate multi-term equations involving integer parameters from [0,9] and elementary functions of the variable $x$, selected from $f(x)\in\{x^{0},x^{1},x^{2},\sin(x),\frac{1}{x},|x|,-x\}$. The agent is provided with the number of functions, the presence of specific parameters in each function (the exact numerical values of these parameters are unknown and represented by letters), and the types of elementary functions involved in the current puzzle. The goal of the agent is to deduce the values of these parameters. For example, in one puzzle, the ground truth is $\mathbf{F_{1}}(x)=a\sin(x)+b\times\frac{1}{x}$, $\mathbf{F_{2}}(x)=ax^{2}$ where $a=3$ and $b=2$. The agent knows the following information: There are three elementary functions in this puzzle $\{\sin(x),\frac{1}{x},x^{2}\}$, there are two functions $\mathbf{F_{1}}(x)$ and $\mathbf{F_{2}}(x)$, $\mathbf{F_{1}}(x)$ has 2 terms and parameters $a,b$ in it, and $\mathbf{F_{2}}(x)$ has 1 term and one parameter $a$. To solve the puzzle, the agent must interact with the environment through a defined action space: selecting a function and assigning values to $x$, then observing the resultant output. For example, assigning values 1 and 2 to $\mathbf{F_{2}}$ reveals a quadratic increase in output, indicating the presence of $x^{2}$ in $\mathbf{F_{2}}$. Similarly, assigning a value of 1 to $\mathbf{F_{1}}$ results in a floating-point output, rather than an integer, suggesting the inclusion of trigonometric components, confirming that $\sin(x)$ is a component of $\mathbf{F_{1}}$. The difficulty of each puzzle is controlled by variations in the number of functions, unknown parameters, and elementary functions in use. We manually enumerated 100 combinations of functions, incorporating different numbers of terms and types of elementary functions to ensure a diverse range of puzzle complexity (see detailed distribution in Table 5 in Appendix  A.2).
@@
The Escape Room. This environment simulates scenarios where no established knowledge is applicable, challenging agents to rely on basic human priors—such as counting, mapping, and attribute abstraction—to convert qualitative observations into general rules through iterative feedback. We create a fictitious setting: an agent is placed in an art gallery escape room and must decipher a 3-digit password to unlock a code-secured door. Each digit of the password represents the count of paintings of a specific type—watercolor, oil, or acrylic—that share a given color. The agent receives brief descriptions of paintings, such as “This is an acrylic painting of a green jungle”, indicating their type and color. Initially, the agent only knows the password is a 3-digit number and is given a hint about which color to focus on. After proposing a hypothesis and entering a password guess, the door provides feedback on which digits are correct, allowing the agent to refine its hypothesis. To prevent a brute-force approach, the specific color associated with the password changes after every three failed attempts. Each puzzle varies in the number of paintings, and while paintings in the same room are visible, those in other rooms remain hidden until the agent moves to access them. The difficulty of this puzzle type is controlled by the different number of paintings, whether agent need to as detailed in Table 6 in the Appendix  A.2.
@@
The Reactor. This environment simulates scenarios without pre-established knowledge, requiring agents to perform sequential, interdependent actions to uncover ordering and transformation patterns—mirroring real-world experimental design, where each step influences the next. Specifically, the agent’s task is to synthesize target strings using a reactor governed by a hidden string-combining rule. These strings are represented by sequences of alphabetic letters, such as A, B, AABB, and CAB. The reactor permits the agent to input two strings, initiating a reaction that produces a new string for use in subsequent experiments. The agent’s objective is to decipher the specific rules that govern string synthesis by methodically testing different string combinations, with the ultimate goal of synthesizing the target string using the discovered rules. We have designed four types of rules, detailed in Table 1. The difficulty of this puzzle type is controlled by the specific rules used, the length of the target string to be synthesized and the number of unique letters contained in the target string, as detailed in Table 7 in Appendix A.2.
@@
Together, these puzzle types simulate a broad spectrum of real-world rule learning by requiring agents to apply both formal knowledge and commonsense reasoning. Detailed statistics for each puzzle type and example puzzles are provided in Appendix A.2 and A.7. The RULEARN benchmark is designed to emulate realistic, complex text environments with diverse rules. To preserve this realism, we do not restrict rule representations to a specific formal language; instead, LLM agents use natural language to describe rules, promoting generalizability and preventing prior knowledge that could undermine the challenge.
@@ 4 The IDEA Agent
The IDEA agent begins with an abductive action to generate an initial hypothesis, followed by a deduction step to create a new plan. Based on this plan, the agent interacts with the environment. Upon receiving feedback from the environment as a new observation, the agent may take an inductive action to refine the hypothesis or perform another interaction with the environment. Deductive action is invoked to adjust the plan every time the hypothesis changes. This reasoning loop continues until the puzzle is solved or a maximum number of steps is reached. After each step, the results are appended to the agent’s memory, including interaction outcomes and any modifications to the hypothesis or plan. We provide a simplified algorithm demonstrating how the IDEA agent operates in Algorithm 1. Specifically, at each step, we prompt the LLM to reflect on the information recorded in the IDEA agent’s components to make decisions and take actions. Detailed prompts for each type of action are available in Appendix A.5.
More detailed implementation of the agent can be found in Appendix A.3.1. Similar to real-life scenarios, when agents solve tasks in RULEARN puzzles, they do not know the outcomes in advance. Consequently, it is challenging to decide when to refine or change their hypothesis and plans, as well as what interactive actions to take to gather pattern-revealing observations. A detailed example of the IDEA agent solving the Reactor Puzzle is provided in Figure 3.